The following observations have been arranged in ascending order. If the median of the data is 63, find the value of x.
29, 32, 48, 50, x, % + 2, 72, 78, 84, 95;

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Given observations: 29, 32, 48, 50, x, x + 2, 72, 78, 84, 95. Here n = 10, Median is mean of the values of the observations at 1 ;th, and ( 1 + 1) th i.e., 5th and 6th 2 www.cbsepdf.com Si1'TISTICS 229 places. :. Median = x + (.x + 2) => 2.r + 2 = 63 2 2 => 2.r = 126 - 2 = 124 => x = 62.